# TrackID3x3: A Dataset and Algorithm for Multi-Player Tracking with Identification and Pose Estimation in 3x3 Basketball Full-court Videos

### v1

@@
Abstract
Multi-object tracking, player identification, and pose estimation are fundamental components of sports analytics, essential for analyzing player movements, performance, and tactical strategies. However, existing datasets and methodologies primarily target mainstream team sports such as soccer and conventional 5-on-5 basketball, often overlooking scenarios involving fixed-camera setups commonly used at amateur levels, less mainstream sports, or datasets that explicitly incorporate pose annotations. In this paper, we propose the TrackID3x3 dataset, the first publicly available comprehensive dataset specifically designed for multi-player tracking, player identification, and pose estimation in 3x3 basketball scenarios. The dataset comprises three distinct subsets (Indoor fixed-camera, Outdoor fixed-camera, and Drone camera footage), capturing diverse full-court camera perspectives and environments. We also introduce the Track-ID task, a simplified variant of the game state reconstruction task that excludes field detection and focuses exclusively on fixed-camera scenarios. To evaluate performance, we propose a baseline algorithm called Track-ID algorithm, tailored to assess tracking and identification quality. Furthermore, our benchmark experiments, utilizing recent multi-object tracking algorithms (e.g., BoT-SORT-ReID) and top-down pose estimation methods (HRNet, RTMPose, and SwinPose), demonstrate robust results and highlight remaining challenges. Our dataset and evaluation benchmarks provide a solid foundation for advancing automated analytics in 3x3 basketball. Dataset and code will be available at https://github.com/open-starlab/TrackID3x3.
Figure 1: Examples of TrackID3x3 dataset with bounding boxes and pose (10 keypoints). We provide indoor fixed-camera (top: Indoor), outdoor fixed-camera (middle: Outdoor), outdoor drone-camera (bottom: Drone) videos in 3x3 basketball. We estimate the player coordinates on the court in the Indoor and Outdoor datasets.
@@ Abstract
SN-Tracking [13] 225,375 Soccer $\times$ outdoor broadcast camera
DeepSport-Basketball [25] 672 Basketball $\checkmark$ indoor broadcast camera
DanceTrack [74] 105,855 Dance $\times$ indoor/outdoor various cameras
@@
SportsMOT [14] 1879 Soccer/Basketball/Volleyball $\times$ indoor/outdoor broadcast camera
SoccerTrack[64] 82,800 Soccer $\times$ outdoor fisheye/drone
@@
TeamTrack [65] 279,900 Soccer/Basketball/Handball $\times$ indoor/outdoor fisheye/drone
SN-GSR [70] 150,000 Soccer $\times$ outdoor broadcast camera
@@
TrackID3x3 (ours) 155,797 3x3 Basketball $\checkmark$ indoor/outdoor smartphone/consumer/drone
Table 1: Comparative overview of MOT and related datasets in sports.
1 Introduction
@@
Recent advances in computer vision have developed sports analytics by enabling the automated extraction of critical information from video data, reducing manual intervention and enhancing analytical precision. Researchers have developed techniques for field registration (e.g., [77, 26, 57]), multi-object tracking (MOT, e.g., [93, 28, 72]), re-identification (Re-ID, e.g., [54, 68, 69]), and pose estimation (e.g., [87, 88, 10]). These approaches also address common sports-specific challenges, such as frequent occlusions and the inherent similarity in visual features among players, which complicate accurate data extraction. These will also support applications ranging from automated decision making and scoring to comprehensive performance analytics by converting raw broadcast footage into structured data. However, the current research is predominantly focused on sports such as soccer, 5-on-5 basketball, and volleyball, while many sports—including 3x3 basketball—lack dedicated datasets and tailored methodologies.
@@
Beyond soccer, the availability of tracking data enough for machine learning analytics in other team sports remains critically limited (note that, the datasets in sports computer vision areas are usually not sufficient for data analytics using machine learning). For instance, the only large-scale tracking dataset available for 5-on-5 basketball is derived from the 2015–16 NBA season, including 630 games; however, it has been outdated now and hinder further research progress. In contrast, soccer game state reconstruction [70]—which integrates field detection, player tracking, and identification in broadcast videos—has been pioneering and remains essential for subsequent analyses such as event prediction [16, 78, 92, 91], trajectory prediction [76, 47, 22], and reinforcement learning [55, 21]. To enable such analytical techniques across sports, it is necessary to develop methods that can effectively process video footage captured by simple, fixed cameras (e.g., smartphones). Establishing robust image processing approaches for these conditions is expected to democratize automated analyses based on machine learning, including prediction with mathematical modeling [71, 40, 34], evaluation [80, 17, 23], and suggestion of optimal actions [89, 55]. In particular, 3x3 basketball offers a promising research opportunity: its reduced number of players simplifies occlusion problems and streamlines problem formulation in computer vision. However, a significant challenge remains in the absence of annotated datasets for player positions and pose estimation in this sport.
@@
In this paper, we propose TrackID3x3 dataset and algorithm for multi-player tracking with identification and pose estimation in 3x3 basketball full-court videos (see Figure 1 and Table 1). Our dataset, the first publicly available resource specifically for 3x3 basketball, includes subsets captured from Indoor (fixed camera), Outdoor (fixed camera), and Drone (drone camera) scenarios, providing diversity in terms of camera perspectives and playing environments. We also propose a new task termed Track-ID task, a simplified version of the game state reconstruction (GSR) task [70], excluding field detection to specifically target fixed-camera scenarios, thereby enabling straightforward subsequent data analysis. Additionally, we establish baseline algorithms using a combination of recent tracking and identification methods. Our benchmark evaluation incorporates recent algorithms for MOT and 2D pose estimation.
@@
The contributions of this paper are as follows. (1) We propose a new comprehensive annotation dataset in 3x3 basketball full-court videos captured under diverse conditions—including indoor and outdoor with smartphone, drone, and consumer camera recordings—addressing the current lack of publicly available resources for this sport (Table 1). (2) We examined the feasibility of fully automating multi-player tracking with identification—a simplified variant of game state reconstruction [70]—in 3x3 basketball, a sport that offers more favorable conditions compared to soccer. Our experimental results demonstrate the effectiveness of our baseline approach, achieving 80.75% in the Indoor dataset and 46.11% in the Outdoor dataset for the proposed TI-HOTA metric (described in Section 4). (3) We present and validate them as the pose dataset, an essential component for future subsequent sports analytics. We establish benchmarks for pose estimation on our dataset by evaluating a range of models from conventional methods to recent top-down transformer-based models, reporting over 77% in all datasets in terms of mean Percentage of Detected Joints (PDJ).
@@
Drone dataset 4,999 29,994 500 Outdoor Drone (DJI Air 2S) 3840 × 2160
@@
Table 2: Overview of our TrackID3x3 Dataset. The unit of pose is frames and that of resolutions is pixel.
2 Related Work
@@
Multi-player tracking in sports. In recent years, significant progress has been made in building datasets. Early efforts provided soccer player location data using multi-view camera [19], while subsequent datasets combined high-resolution panorama videos with local positioning system data [60]. Publicly available large-scale broadcast video datasets such as SoccerNet [13, 70] and SoccerDB [36] have further propelled research in player tracking. In basketball, datasets including APIDIS and SPIROUDOME [15, 49] have been introduced, alongside other sports datasets for handball [8] and volleyball [31]. Recent datasets such as SoccerNet-Tracking [13] and SportsMOT [14] utilize unedited broadcast footage, while SoccerTrack [64] and its extension TeamTrack [65] offer full-pitch multi-sport tracking. Furthermore, virtual environments like Google Research Football [44] and Soccersynth [62] provide synthetic data that facilitate algorithm benchmarking.
@@
In recent MOT, tracking-by-detection is a popular approach, where high-quality object detection forms the backbone for subsequent data association and tracking. The first step is the accurate identification of players and the ball in each video frame. Early methods relied on background subtraction, multi-scale feature detection, and edge detection techniques [51, 11, 18, 3]. More recently, deep learning–based detectors such as RetinaNet [46], CenterNet [20], and various versions of YOLO [63] have significantly improved detection accuracy. The second step associates detections across consecutive frames, which is challenging in sports due to frequent occlusions, rapid motion, and visual similarity among players. Motion-based methods, such as SORT [7] and its extensions like BoT-SORT [1], employ Kalman filters and data association techniques to predict object trajectories. Deep appearance-based methods, exemplified by DeepSORT [85] and techniques leveraging Deep EIoU [30], use convolutional neural networks to extract robust visual features that help maintain consistent identities despite occlusion and similar uniforms. Recent work in soccer [52, 53], basketball [5, 28, 93], hockey [83], and multi-sports [29], combines these strategies to tackle the unique challenges of sports tracking, including complex movements, similar appearances, and severe occlusions. In tracking and dataset diversity, 3x3 basketball reduces the number of players simplifies occlusion and other problems.
@@
Player identification in sports. Player identification, including Re-ID (the process of consistently matching and maintaining a player’s identity across different frames and camera views) is fundamental to sports analytics. In non-team sports, marathon, ultra running, and running datasets [6, 58, 59, 75] provide images or videos of athletes annotated with or without bib numbers. In team sports, re-ID datasets have been proposed in basketball [81], ice hockey [41, 96], and soccer [12]. Recently, the SoccerNet GSR dataset [70] provided a comprehensive dataset integrating field registration, tracking, and player identification.
@@
Player Re-ID methods have evolved from relying solely on traditional visual cues such as jersey color and number [33, 56]. Early image-level methods focused on facial and local appearance features, while more recent work employs convolutional neural networks and part-based methods [66, 2]. Advanced techniques now incorporate contrastive learning and transformer-based models [42, 82, 27], as well as multi-task frameworks for joint Re-ID, team affiliation, and role classification (PRTreID) [54]. In GSR [70] baseline, player identification was performed with PRTreID [54] to produce team and role-aware ReID embeddings, and MMOCR [43] for jersey number recognition. We use a general framework [41] for jersey number recognition and color histograms for tracklet integration and team affiliation determination for both jersey and bibs situations.
Pose estimation in sports. Pose information is fundamental in sports analytics because it provides detailed insights into players’ movements, enabling precise performance assessments. In general sports contexts, LSP dataset offers 1,000 images with pose [38], and the Sport Image dataset with 1,300 images [84]. The extensive MPII Human Pose dataset provides 25,000 images covering a wide range of human motions, including sports scenarios [4], while the OCHuman dataset emphasizes occluded human poses with 4,700 images [94]. For comprehensive whole-body pose annotations, the COCO-WholeBody dataset [37] has been developed, and domain-specific datasets such as 3DSP [90] and the KTH Multiview Football Dataset II [39] provide valuable pose data for soccer players. In addition, the DeepSportLab dataset [25] offers a unified resource for ball detection, player instance segmentation, and pose estimation (672 frames) in basketball contexts. Our dataset provides 6701 frames of annotated pose (Table 2).
@@
Recent advances in 2D pose estimation have been primarily driven by top-down methods in sports (e.g., [90, 93]), where the process begins with detecting individual persons followed by keypoint localization. The top-down paradigm consistently yields higher precision in keypoint estimation in sports scenes characterized by frequent occlusions and rapid motion. DeepPose [79] pioneered this approach, and subsequent work such as HRNet [73] further improved accuracy by preserving high-resolution representations throughout the network. More recent top-down systems, including RTMPose [35] and SwinPose [87], leverage transformer-based architectures to further refine keypoint predictions. Although bottom-up methods like DeepCut [61] and OpenPose [9] as well as emerging end-to-end approaches [67, 48] provide efficient alternatives, we focus our evaluation solely on top-down methods because bounding boxes are given in our datasets.
3 TrackID3x3 Dataset

### v2

@@
Abstract.
Multi-object tracking, player identification, and pose estimation are fundamental components of sports analytics, essential for analyzing player movements, performance, and tactical strategies. However, existing datasets and methodologies primarily target mainstream team sports such as soccer and conventional 5-on-5 basketball, often overlooking scenarios involving fixed-camera setups commonly used at amateur levels, less mainstream sports, or datasets that explicitly incorporate pose annotations. In this paper, we propose the TrackID3x3 dataset, the first publicly available comprehensive dataset specifically designed for multi-player tracking, player identification, and pose estimation in 3x3 basketball scenarios. The dataset comprises three distinct subsets (Indoor fixed-camera, Outdoor fixed-camera, and Drone camera footage), capturing diverse full-court camera perspectives and environments. We also introduce the Track-ID task, a simplified variant of the game state reconstruction task that excludes field detection and focuses exclusively on fixed-camera scenarios. To evaluate performance, we propose a baseline algorithm called Track-ID algorithm, tailored to assess tracking and identification quality. Furthermore, our benchmark experiments, utilizing recent multi-object tracking algorithms (e.g., CAMELTrack) and top-down pose estimation methods (HRNet, RTMPose, and SwinPose), demonstrate robust results and highlight remaining challenges. Our dataset and evaluation benchmarks provide a solid foundation for advancing automated analytics in 3x3 basketball. Dataset and code will be available at https://github.com/open-starlab/TrackID3x3.
Dataset, Sports, Multi-object tracking, Human pose estimation
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 8th International ACM Workshop on Multimedia Content Analysis in Sports; October 27–28, 2025; Dublin, Irelandbooktitle: Proceedings of the 8th International ACM Workshop on Multimedia Content Analysis in Sports (MMSports ’25), October 27–28, 2025, Dublin, Irelanddoi: 10.1145/3728423.3759400isbn: 979-8-4007-1835-9/2025/10ccs: Computing methodologies Computer vision tasks Activity recognition and understandingccs: Computer systems organization Architectures Other architectures Special purpose systems
Figure 1. Examples of TrackID3x3 dataset with bounding boxes and pose (10 keypoints). We provide indoor fixed-camera (top: Indoor), outdoor fixed-camera (middle: Outdoor), outdoor drone-camera (bottom: Drone) videos in 3x3 basketball. We estimate the player coordinates on the court in the Indoor and Outdoor datasets.
@@ Abstract.
SN-Tracking (Cioppa et al., 2022b) 225,375 Soccer $\times$ outdoor broadcast camera
DeepSport-Basketball (Ghasemzadeh et al., 2021) 672 Basketball $\checkmark$ indoor broadcast camera
DanceTrack (Sun et al., 2022) 105,855 Dance $\times$ indoor/outdoor various cameras
@@
SportsMOT (Cui et al., 2023) 150,379 Soccer/Basketball/Volleyball $\times$ indoor/outdoor broadcast camera
SoccerTrack(Scott et al., 2022) 82,800 Soccer $\times$ outdoor fisheye/drone
@@
TeamTrack (Scott et al., 2024) 279,900 Soccer/Basketball/Handball $\times$ indoor/outdoor fisheye/drone
SN-GSR (Somers et al., 2024b) 150,000 Soccer $\times$ outdoor broadcast camera
@@
TrackID3x3 (ours) 277,926 3x3 Basketball $\checkmark$ indoor/outdoor smartphone/consumer/drone
Table 1. Comparative overview of MOT and related datasets in sports.
1. Introduction
@@
Recent advances in computer vision have developed sports analytics by enabling the automated extraction of critical information from video data, reducing manual intervention and enhancing analytical precision. Researchers have developed techniques for field registration (e.g., (Theiner and Ewerth, 2023; Gutiérrez-Pérez and Agudo, 2024)), multi-object tracking (MOT, e.g., (Yin et al., 2024; Hu et al., 2024; Sun et al., 2024)), re-identification (Re-ID, e.g., (Mansourian et al., 2023; Somers et al., 2023, 2024a)), and pose estimation (e.g., (Xiong et al., 2022; Xu et al., 2022; Cheng et al., 2020)). These approaches also address common sports-specific challenges, such as frequent occlusions and the inherent similarity in visual features among players, which complicate accurate data extraction. These will also support applications ranging from automated decision making and scoring to comprehensive performance analytics by converting raw broadcast footage into structured data. However, the current research is predominantly focused on sports such as soccer, 5-on-5 basketball, and volleyball, while many sports—including 3x3 basketball—lack dedicated datasets and tailored methodologies.
@@
Beyond soccer, the availability of tracking data enough for machine learning analytics in other team sports remains critically limited (note that, the datasets in sports computer vision areas are usually not sufficient for data analytics using machine learning). For instance, the only large-scale tracking dataset available for 5-on-5 basketball is derived from the 2015–16 NBA season, including 630 games; however, it has been outdated now and hinder further research progress. In contrast, soccer game state reconstruction (Somers et al., 2024b)—which integrates field detection, player tracking, and identification in broadcast videos—has been pioneering and remains essential for subsequent analyses such as event prediction (Decroos et al., 2019; Toda et al., 2022; Yeung et al., 2025b, a), trajectory prediction (Lindström et al., 2020; Fujii et al., 2024a), and reinforcement learning (Nakahara et al., 2023; Fujii et al., 2023). To enable such analytical techniques across sports, it is necessary to develop methods that can effectively process video footage captured by simple, fixed cameras (e.g., smartphones). Establishing robust image processing approaches for these conditions is expected to democratize automated analyses based on machine learning, including prediction with mathematical modeling (Spearman, 2018; Kono and Fujii, 2024; Iwashita et al., 2024), evaluation (Ding et al., 2024; Fujii et al., 2024b), and suggestion of optimal actions (Yeung and Fujii, 2024; Nakahara et al., 2023). In particular, 3x3 basketball offers a promising research opportunity: its reduced number of players simplifies occlusion problems and streamlines problem formulation in computer vision. However, a significant challenge remains in the absence of annotated datasets for player positions and pose estimation in this sport.
@@
In this paper, we propose TrackID3x3 dataset and algorithm for multi-player tracking with identification and pose estimation in 3x3 basketball full-court videos (see Figure 1 and Table 1). Our dataset, the first publicly available resource specifically for 3x3 basketball, includes subsets captured from Indoor (fixed camera), Outdoor (fixed camera), and Drone (drone camera) scenarios, providing diversity in terms of camera perspectives and playing environments. We also propose a new task termed Track-ID task, a simplified version of the game state reconstruction (GSR) task (Somers et al., 2024b), excluding field detection to target fixed-camera scenarios, thereby enabling straightforward processes (see Section 4.1). Additionally, we establish baseline algorithms using a combination of recent tracking and identification methods. Our benchmark evaluation incorporates recent algorithms for MOT and 2D pose estimation.
@@
The contributions of this paper are as follows. (1) We propose a new comprehensive annotation dataset in 3x3 basketball full-court videos captured under diverse conditions—including indoor and outdoor with smartphone, drone, and consumer camera recordings—addressing the current lack of publicly available resources for this sport (Table 1). (2) We examined the feasibility of fully automating multi-player tracking with identification—a simplified variant of game state reconstruction (Somers et al., 2024b)—in 3x3 basketball, a sport that offers more favorable conditions compared to soccer. Our experimental results demonstrate the effectiveness of our baseline approach, achieving 85.53% in the Indoor dataset and 71.03% in the Outdoor dataset for the proposed TI-HOTA metric (described in Section 4). (3) We present and validate them as the pose dataset, an essential component for future subsequent sports analytics. We establish benchmarks for pose estimation on our dataset by evaluating a range of models from conventional methods to recent top-down transformer-based models, reporting over 77% in all datasets in terms of mean Percentage of Detected Joints (PDJ).
@@
Drone dataset 127,119 762,714 500 outdoor Drone (DJI Air 2S) 3840 × 2160
@@
Table 2. Overview of our TrackID3x3 Dataset. The unit of pose is frames and that of resolutions is pixel.
2. Related Work
@@
Multi-player tracking in sports. In recent years, significant progress has been made in building datasets. Early efforts provided soccer player location data using multi-view camera (D’Orazio et al., 2009), while subsequent datasets combined high-resolution panorama videos with local positioning system data (Pettersen et al., 2014). Publicly available large-scale broadcast video datasets such as SoccerNet (Cioppa et al., 2022b; Somers et al., 2024b) and SoccerDB (Jiang et al., 2020) have further propelled research in player tracking. In basketball, datasets including APIDIS and SPIROUDOME (De Vleeschouwer et al., 2008; Lu et al., 2017) have been introduced, alongside other sports datasets for handball (Biermann et al., 2021) and volleyball (Ibrahim et al., 2016). Recent datasets such as SoccerNet-Tracking (Cioppa et al., 2022b) and SportsMOT (Cui et al., 2023) utilize unedited broadcast footage, while SoccerTrack (Scott et al., 2022) and its extension TeamTrack (Scott et al., 2024) offer full-pitch multi-sport tracking. Furthermore, virtual environments like Google Research Football (Kurach et al., 2020) and Soccersynth (Qin et al., 2025) provide synthetic data that facilitate algorithm benchmarking.
@@
In recent MOT, tracking‐by‐detection is a popular approach, where high-quality object detection forms the backbone for subsequent data association and tracking. The first step is the accurate identification of players and the ball in each video frame. Early methods relied on background subtraction, multi-scale feature detection, and edge detection techniques (Mackowiak et al., 2010; Cheshire et al., 2015; Direkoglu et al., 2018). More recently, deep learning–based detectors such as RetinaNet (Lin et al., 2017), CenterNet (Duan et al., 2019), and various versions of YOLO (Redmon et al., 2016) have significantly improved detection accuracy. The second step associates detections across consecutive frames, which is challenging in sports due to frequent occlusions, rapid motion, and visual similarity among players. Motion-based methods, such as SORT (Bewley et al., 2016) and its extensions like BoT-SORT (Aharon et al., 2022), employ Kalman filters and data association techniques to predict object trajectories. Deep appearance-based methods, exemplified by DeepSORT (Wojke et al., 2017) and techniques leveraging Deep-EIoU (Huang et al., 2024), use convolutional neural networks to extract robust visual features that help maintain consistent identities despite occlusion and similar uniforms. In addition, the recently proposed CAMELTrack (Somers et al., 2025) incorporates an association method based on a learnable module rather than the heuristic associations like the above methods. Recent works in soccer (Maglo et al., 2022; Majeed et al., 2024), basketball (Hu et al., 2024; Yin et al., 2024), hockey (Vats et al., 2023), and multi-sports (Huang et al., 2023; Sun et al., 2024; Khanna et al., 2025; Stanczyk et al., 2025), combine these strategies and their own unique methods to tackle the unique challenges of sports tracking, including complex movements, similar appearances, and severe occlusions. The proposed dataset provides experimental settings that simplify issues in basketball player tracking, such as multi-player occlusion (Hu et al., 2024), as well as camera positions and bibs that are not found in existing benchmark datasets.
@@
Player identification in sports. Player identification, including Re-ID (the process of consistently matching and maintaining a player’s identity across different frames and camera views) is fundamental to sports analytics. In non-team sports, marathon, ultra running, and running datasets (Ben-Ami et al., 2012; Penate-Sanchez et al., 2020; Suzuki et al., 2024) provide images or videos of athletes annotated with or without bib numbers. In team sports, re-ID datasets have been proposed in basketball (Van Zandycke et al., 2022), ice hockey (Koshkina and Elder, 2024; Zhao et al., 2020), and soccer (Cioppa et al., 2022a). Recently, the SoccerNet GSR dataset (Somers et al., 2024b) provided a comprehensive dataset integrating field registration, tracking, and player identification.
@@
Player Re-ID methods have evolved from relying solely on traditional visual cues such as jersey color and number (Ivankovic et al., 2014). Early image-level methods focused on facial and local appearance features, while more recent work employs convolutional neural networks and part-based methods (Akan and Varlı, 2023). Advanced techniques now incorporate contrastive learning and transformer-based models (Koshkina et al., 2021; Vats et al., 2022), as well as multi-task frameworks for joint Re-ID, team affiliation, and role classification (PRTreID) (Mansourian et al., 2023). In GSR (Somers et al., 2024b) baseline, player identification was performed with PRTreID (Mansourian et al., 2023) to produce team and role-aware ReID embeddings, and MMOCR (Kuang et al., 2021) for jersey number recognition. We use a general framework (Koshkina and Elder, 2024) for jersey number recognition and color histograms, and relative position on the court to determine team affiliation.
Pose estimation in sports. Pose information is fundamental in sports analytics, offering detailed insights into players’ movements for precise performance assessment. In MOT, pose annotations provide fine-grained structural cues for Re-ID or tracklet association. In scenarios with occlusions or similar uniforms, keypoint-based features help reduce ID switches. Many studies have leveraged pose information to improve MOT (Iqbal et al., 2017; Yin et al., 2024). While our pose annotations are not solely for training MOT baselines, they offer multimodal supervision (bounding box + pose) to support future research and pose-assisted MOT approaches.
In sports contexts, the LSP dataset contains 1,000 pose-labeled images (Johnson and Everingham, 2010), the Sport Image dataset has 1,300 (Wang et al., 2011), and COCO-WholeBody provides comprehensive whole-body annotations (Jin et al., 2020). Domain-specific datasets include 3DSP (Yeung et al., 2024) and DeepSportLab for basketball, which offers ball detection, player segmentation, and pose estimation (672 frames) (Ghasemzadeh et al., 2021). Our pose annotations aim to provide sufficient samples for method development and validation, covering diverse scenes, viewpoints, and occlusions. Based on our experience, the 6,701 annotated frames are sufficient for research and evaluation (Table 2).
@@
Recent advances in 2D pose estimation have been primarily driven by top-down methods in sports (e.g., (Yeung et al., 2024; Yin et al., 2024)), where the process begins with detecting individual persons followed by keypoint localization. The top-down paradigm consistently yields higher precision in keypoint estimation in sports scenes characterized by frequent occlusions and rapid motion. DeepPose (Toshev and Szegedy, 2014) pioneered this approach, and subsequent work such as HRNet (Sun et al., 2019) further improved accuracy by preserving high-resolution representations throughout the network. More recent top-down systems, including RTMPose (Jiang et al., 2023) and SwinPose (Xiong et al., 2022), leverage transformer-based architectures to further refine keypoint predictions. Although bottom-up methods like DeepCut (Pishchulin et al., 2016) and OpenPose (Cao et al., 2018) as well as emerging end-to-end approaches (Shi et al., 2022; Liu et al., 2023) provide efficient alternatives, we focus our evaluation solely on top-down methods because bounding boxes are given in our datasets.
3. TrackID3x3 Dataset